# Information-Consistent Language Model Recommendations through Group Relative Policy Optimization

### v1

@@ 5.2 GRPO Training Setup
The experiments used Unsloth’s unsloth GRPO implementation applied to the Llama-3.2-1B-Instruct meta2024llama3.2-1b-instruct; unsloth2024llama3.2-1b-instruct model with LoRA adapters for parameter-efficient fine-tuning. The GRPO trainer was configured with the combined reward function described in Section IV-B. The configuration parameters were as follows:
@@
  • Dataset: Benchmark dataset from prabhune2025llms, with repeated male/female prompt pairs
@@ 6 RESULTS
We evaluate our approach using the RealWorldQuestioning Benchmark dataset introduced in prabhune2025llms, designed to measure gender-related differences in information content in LLM outputs. We chose this dataset because it has real-world queries which are similar in every aspect except for gender differences. This structure enables systematic evaluation of consistency in entropy of LLM responses under realistic user scenarios and supports robust comparison of LLM behavior across gendered prompt variants.
@@ 7 DISCUSSION
At the same time, several limitations must be acknowledged. First, our evaluation was restricted to a controlled set of gender-based prompt variations in the investment/job recommendation domain. While this setup offers a clear demonstration of inconsistency, it does not cover the full spectrum of prompt variability seen in real-world use. Second, context was deliberately kept at zero in our experiments (fresh conversations for every prompt) to ensure equivalency. Although this isolates prompt effects, many enterprise deployments involve multi-turn dialogue with accumulated context, where consistency requirements may differ. Third, entropy-based metrics, while useful as proxies for informativeness and stability, may not fully capture the qualitative dimensions of consistency valued by end users. Finally, our experiments were conducted on the 1B parameter Llama-3.2 instruct model; results may vary for larger or differently tuned architectures.
These limitations point toward important directions for future work. Expanding beyond gender to include paraphrasing, tone, or regional variations will help validate generalizability. Evaluating consistency in multi-turn conversations will clarify how stability objectives interact with personalization. Additionally, exploring richer evaluation metrics—such as semantic similarity, factual overlap, or user satisfaction studies—could provide a more holistic assessment of consistency in enterprise contexts.

### v2

@@ 5.2 GRPO Training Setup
The experiments used Unsloth’s Daniel Han and team (2023) GRPO implementation applied to the Llama-3.2-1B-Instruct AI (2024a, b) model with LoRA adapters for parameter-efficient fine-tuning. The GRPO trainer was configured with the combined reward function described in Section IV-B. The configuration parameters were as follows:
@@
  • Dataset: Benchmark dataset from Prabhune et al. (2025), with repeated male/female prompt pairs
@@ 6 RESULTS
We evaluate our approach using the RealWorldQuestioning Benchmark dataset introduced in Prabhune et al. (2025), designed to measure gender-related differences in information content in LLM outputs. We chose this dataset because it has real-world queries which are similar in every aspect except for gender differences. This structure enables systematic evaluation of consistency in entropy of LLM responses under realistic user scenarios and supports robust comparison of LLM behavior across gendered prompt variants.
@@ 7 DISCUSSION
At the same time, several limitations must be acknowledged. First, our evaluation was restricted to a controlled set of gender-based prompt variations in the investment/job recommendation domain. While this setup offers a clear demonstration of inconsistency, it does not cover the full spectrum of prompt variability seen in real-world use. Second, context was deliberately kept at zero in our experiments (fresh conversations for every prompt) to ensure equivalency. Although this isolates prompt effects, many enterprise deployments involve multi-turn dialogue with accumulated context, where consistency requirements may differ. Third, entropy-based metrics, while useful as proxies for informativeness and stability, may not fully capture the qualitative dimensions of consistency valued by end users. Finally, our experiments were conducted on the 1B parameter Llama-3.2 instruct model; results may vary for larger or differently tuned architectures. These limitations point toward important directions for future work. Expanding beyond gender to include paraphrasing, tone, or regional variations will help validate generalizability. Evaluating consistency in multi-turn conversations will clarify how stability objectives interact with personalization. Additionally, exploring richer evaluation metrics—such as semantic similarity, factual overlap, or user satisfaction studies—could provide a more holistic assessment of consistency in enterprise contexts.